# RoboMatch: A Mobile-Manipulation Teleoperation Platform with Auto-Matching Network Architecture for Long-Horizon Manipulation

### v1

@@ I INTRODUCTION
Imitation learning has gained significant attention in robotics due to its practical applicability. Strategies like ACT [5] and Diffusion Policy [6] enable autonomous decision-making through end-to-end learning from demonstrations and environmental perception. Such perception includes visual [7] and proprioceptive sensing [8]. While visual representations encode scene information, learning robust visual features for robotics remains challenging. Attention mechanisms like the Efficient Multi-scale Attention Module (EMA) [9] enhance visual perception but focus only on temporal features, ignoring frequency-domain information. Proprioceptive accuracy critically affects motion control precision. During delicate tasks like assembly or grasping, insufficient data and error accumulation often cause end-effector pose deviations, leading to failure. Improving proprioception and reducing error propagation are thus key challenges in embodied intelligence.
@@
2) Proprioceptive-Visual Enhanced Diffusion Policy: We design the PVE-DP algorithmic framework, achieving breakthroughs in two key aspects:
We integrate the FE-EMA visual module (Frequency Enhanced EMA) into the Diffusion Policy visual backbone network to generate spatio-frequency fused visual representations.
We embed IMU sensors on both end-effectors to accurately perceive rotational states by capturing real-time quaternion data from each arm. This information is fused with joint angles to form a comprehensive proprioceptive representation.
@@ II-B Imitation Learning
Despite their strong task adaptability, diffusion models still exhibit noticeable limitations in motion generation precision, especially in fine manipulation tasks (e.g., folding clothes) that rely solely on joint angles and simple temporal image observations. Cumulative errors during inference, coupled with insufficient visual perception and proprioception, significantly impede precise decision-making and efficient robot control [36].
@@ III-B The Model Framework of PVE-DP
As shown in Fig. 3, to enable the robot to capture richer perceptual data, we propose the Proprioceptive-Visual Enhanced Diffusion Policy (PVE-DP). Building upon the original Diffusion Policy, this approach incorporates a proprioceptive enhancement module and a visual enhancement module. First, we integrate IMUs at the robot arm’s end-effectors to acquire quaternion data , which satisfies the unit quaternion constraint:
The real-time end-effector quaternion is incorporated as a new modality and further fused with joint angle data to construct a rich proprioceptive feature representation
At the algorithmic level, we introduce a Frequency Enhanced EMA (FE-EMA) module and deeply integrate it with the visual backbone of the diffusion policy. This module enables efficient fusion of temporal and frequency domain features, thereby constructing time-frequency visual feature representations.
Given an input feature map , its wavelet transform is expressed by Equation (4), achieved by constructing a set of scaled and translated wavelet basis functions and scaling functions :
@@
The high-frequency feature fusion is expressed as , combined with the low-frequency feature to form an enhanced feature representation . Finally, the FE-EMA module generates a cross-space attention map via matrix dot-product to capture pixel-wise relationships and enhance feature representation. The temporal fusion feature is generated through an attention mechanism and integrated with the frequency-domain feature , as shown in Equation (5):
In Equation (5), represents the temporal domain feature generated through an attention mechanism and group normalization process. It primarily captures local regional information extracted from the input features. This feature is combined with the frequency-domain feature to ensure that the model captures more comprehensive multi-scale information. Such an approach enables dynamic balancing between temporal and frequency domain features, effectively enhancing the model’s performance.
Based on the above multi-modal fusion representation, the loss function of the diffusion policy is defined as follows:
@@
During the denoising process, PVE-DP continuously refines the predicted action sequence by incorporating additional wrist perception data from robotic arms IMUs, along with time-frequency features extracted by the FE-EMA network, enabling precise motion execution in fine manipulation tasks.
@@ III-C Auto-Matching Network (AMN) Architecture
As shown in Fig. 4, we propose the Auto-Matching Network (AMN) framework. This architecture integrates the vision-language model GLM-4.1V [12] and employs a “Chain of Thought” reasoning paradigm to decompose complex long-horizon tasks into logically connected subtask sequences. It dynamically matches each subtask with a pre-trained policy network for distributed inference. The core modeling concept can be expressed as:
The framework leverages the Chain of Thought cognitive paradigm to decompose a complex long-horizon task into a sequence of logically connected subtasks . During training, for tasks spanning over 3000+ time steps, we introduce a step-wise saving mechanism: upon completion of each subtask , the current state and action are saved as a subtask sample. Through this temporal decoupling strategy, the original task is decomposed into short-horizon subtasks , each trained with an independent network architecture . Due to the significant differences in perception-action patterns across subtasks, the learned network weights for each subtask exhibit high specificity.
During inference, the operator provides a command via voice or text. The robot observes the scene and sends the image and instruction to the VLM. System decomposes the task into sequential steps, each assigned a pre-trained specialized policy network for execution. The PVE-DP policy is matched to perform robotic manipulation. After each subtask, the system automatically selects the next network until the entire task is completed.

### v2

# RoboMatch: A Unified Mobile-Manipulation Teleoperation Platform with Auto-Matching Network Architecture for Long-Horizon Tasks

@@ I INTRODUCTION
Imitation learning has gained significant attention in robotics due to its practical applicability. Strategies like ACT [5] and Diffusion Policy [6] enable autonomous decision-making through end-to-end learning from demonstrations and environmental perception. Such perception includes visual [7] and proprioceptive sensing [8]. While visual representations encode scene information, learning robust visual features for robotics remains challenging. Attention mechanisms like the Efficient Multi-scale Attention Module (EMA) [9] enhance visual perception but focus only on spatial features, ignoring frequency-domain information. Proprioceptive accuracy critically affects motion control precision. During delicate tasks like assembly or grasping, insufficient data and error accumulation often cause end-effector pose deviations, leading to failure. Improving proprioception and reducing error propagation are thus key challenges in embodied intelligence.
@@
2) Proprioceptive-Visual Enhanced Diffusion Policy: We design the PVE-DP framework, introducing enhancements in two key areas:
We integrate the FE-EMA visual module (Frequency Enhanced EMA) into the Diffusion Policy visual backbone network to construct spatio-frequency-fused visual representations.
We mount IMU sensors on both end-effectors to accurately perceive rotational states by capturing real-time quaternion data from each arm. This information is fused with joint angles to form a comprehensive proprioceptive representation.
@@ II-B Imitation Learning
Despite their strong task adaptability, diffusion models still exhibit noticeable limitations in motion generation precision, especially in fine manipulation tasks (e.g., folding clothes) that rely solely on joint angles and simple spatial image observations. Cumulative errors during inference, coupled with insufficient visual perception and proprioception, significantly impede precise decision-making and efficient robot control [36].
@@ III-B The Model Framework of PVE-DP
As shown in Fig. 3, to enable the robot to capture richer perceptual data, we propose the Proprioceptive-Visual Enhanced Diffusion Policy (PVE-DP). Building upon the original Diffusion Policy, this approach incorporates a proprioceptive enhancement module and a visual enhancement module. First, we integrate IMUs at the robot arm’s end-effectors to acquire quaternion data , which satisfies the unit quaternion constraint in Eq. (1) and Eq. (2):
In Eq. (3), the end-effector quaternion is incorporated as a new modality and further fused with joint angle data to construct a rich proprioceptive feature representation :
At the algorithmic level, we introduce a Frequency Enhanced EMA (FE-EMA) module and deeply integrate it with the visual backbone of the diffusion policy. This module enables efficient fusion of spatial and frequency domain features, thereby constructing spatio-frequency visual feature representations.
Given an input feature map , its wavelet transform is expressed as Eq. (4), achieved by constructing a set of scaled and translated wavelet basis functions and scaling functions :
@@
The high-frequency feature fusion is expressed as , combined with the low-frequency feature to form an enhanced feature representation . Finally, the FE-EMA module generates a cross-space attention map via matrix dot-product to capture pixel-wise relationships and enhance feature representation. The spatial fusion feature is generated through an attention mechanism and integrated with the frequency-domain feature , as shown in Eq. (5):
In Eq. (5), represents the spatial domain feature generated through an attention mechanism and group normalization process. It primarily captures local regional information extracted from the input features. This feature is combined with the frequency-domain feature to ensure that the model captures more comprehensive multi-scale information. Such an approach enables dynamic balancing between spatial and frequency domain features, effectively enhancing the model’s performance.
Based on the above multi-modal fusion representation, the loss function of the diffusion policy is defined in Eq. (6):
@@
During the denoising process, PVE-DP continuously refines the predicted action sequence by incorporating additional wrist perception data from robotic arms IMUs, along with spatio-frequency features extracted by the FE-EMA network, enabling precise motion execution in fine manipulation tasks.
@@ III-C Auto-Matching Network (AMN) Architecture
As shown in Fig. 4, we propose the Auto-Matching Network (AMN) framework. This architecture integrates the vision-language model GLM-4.1V [12] and employs a “Chain of Thought” reasoning paradigm to decompose complex long-horizon tasks into logically connected subtask sequences. It dynamically matches each subtask with a pre-trained policy network for distributed inference. The core modeling concept can be expressed as Eq. (7):
The architecture leverages the Chain of Thought cognitive paradigm to decompose a complex long-horizon task into a sequence of logically connected subtasks . During training, for tasks spanning over 3000+ time steps, we introduce a step-wise saving mechanism: upon completion of each subtask , the current state and action are saved as a subtask sample. Through this temporal decoupling strategy, the original task is decomposed into short-horizon subtasks , the -th of which is trained with an independent network architecture . Due to the significant differences in perception-action patterns across subtasks, the learned network weights for each subtask exhibit high specificity.
During inference, the operator provides a command via voice or text. The robot observes the scene and sends the image and instruction to the VLM. The AMN architecture decomposes the task into sequential steps, and then assigns each step a pre-trained specialized policy network for execution. The PVE-DP policy is matched to perform robotic manipulation. After each subtask, the system automatically selects the next network until the entire task is completed.